OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ...

OpenStudy (anonymous):

21, 30, 39

OpenStudy (anonymous):

nth term= a+ (n-1)d where a=first term d= difference between two terms

OpenStudy (anonymous):

nth term= -15+ (n-1)9 = 9n-24

OpenStudy (anonymous):

is this what you are looking for?

OpenStudy (agent0smith):

@melacho still need help?